when a student subtracts 10 from a number, the results is 1/2 of the number. whats the number?
Answer by Cromlix(4381)

Hi there,
Make 'x' your number:-
x - 10 = x/2
Multiply through by 2
2x - 20 = x
Collect like terms:-
2x - x = 20
x = 10
This is your number.
Hope this helps:-)